# GAMA: High-Performance GEMM Acceleration on AMD Versal ML-Optimized AI Engines

### v2

@@ IV-B Chaining the AIEs together
AIE to AIE connectivity: AIEs are connected in series, as shown in Figure 3, to form a pack, allowing larger GEMMs to be implemented with increased parallelism and reuse. There are two PLIOs for input to each AIE in the pack and one PLIO for output from the pack. Reductions are performed inside the pack, and only final results are written back to the PL. This reduces the output PLIOs in the design, since only the last AIE has to write data back to the PL. Cascade, via-switch connection, and buffer are the various methods for partial sum communication from one AIE to the next AIE. CHARM [20], ARIES [22] and RSN-XNN [15] use a cascade interface, while MaxEVA [14] and AMA [8] use buffer. Buffer communication does not stream data from one AIE to another; it uses buffers in the neighboring AIE’s memory. These buffers consume memory. Via-switch and cascade connections stream data from one AIE to another avoiding this memory overhead. Via-switch interface is only 32-bit and using it increases communication latency significantly. AIE architecture has a 512-bit wide cascade interface compared to AIE1’s 384-bit. The cascade with the larger bit width and no memory overhead becomes the best candidate for AIE to AIE communication. We use cascade interface for all partial sum communication. Cascade connections make the pack run in a data-flow fashion.
@@
Different kernel types: Our design incorporates three distinct kernels within the pack. The first kernel (used in the leftmost AIE in the pack) takes Matrix A and Matrix B as inputs from the PLIO and output’s partial sum through cascade interface. The second kernel (used in the middle AIEs in the pack) has two input PLIO ports, one input cascade port, and one output cascade port. The third kernel (used in the last AIE in the pack) receives two inputs from PLIO, one input from the cascade port, and one output PLIO port.
@@
AIEs connected in a pack perform a GEMM operation with matrix sizes larger than the matrix sizes for a single AIE. For example, if the kernel sizes are M=48, N=240 and N=48 then the native size of the pack becomes M=48, K=960 and N=48.
@@ IV-C Scaling to the complete AIE array
Scaling hyperparameters and constraints: The scaling in our design is governed by three hyperparameters. Y: Replicates the pack along the Y-axis of the AIE array. G: Defines the size of each pack. X: Replicates the pack along the X-axis of the AIE array. Figure 5 shows the scaling for the case (Y = 2,G = 4,X = 2). PLIOs are shown using different colors that show broadcast and reuse. Matrix A and matrix B are assigned to 16 AIE as shown in Figure 5. The same mapping can be scaled to the complete array.
@@
$Y\times X\leq PLIO\_out$ (9)
Kernel placement: Scaling the design to the entire array by satisfying the constraints stated above leads to compilation failure due to PLIO routing congestion. This happens because the third AIE utilizes three distinct PLIOs (two read, one write), leading to congestion in the switch network’s vertical lane when extended across all rows. To avoid this congestion, GAMA uses a zigzag or skewed kernel placement strategy across the array, as shown in Figure 7. We skew the kernel placement by two since skewing by one also leads to congestion, and skewing by three reduces the AIE array utilization. The first two AIEs in each alternate rows of AIEs are not used. This pattern alternates the third AIE’s location in each row, reducing the pressure of each vertical switch lane. This placement successfully avoids congestion, leading to a scalable and compilable design.
@@ V-B Performance of single AIE kernels
At the single AIE level, the two most important aspects are the MMUL API size and the single AIE kernel size (MxKxN). As discussed in Section IV-A, the single AIE kernel size is influenced by multiple factors such as the theoretical compute to communication ratio $\gamma$, AIE memory utilization and the input and output precision.
@@
To calculate the efficiency of the kernel, we measure the kernel compute cycles (KCC) and then take the ratio of theoretical KCC to measured KCC and call it kernel compute efficiency (KCE). We use KCE as a metric to find the best MMUL API size. We sweep all MMUL sizes to find the optimal M, K, and N sizes, as noted in Section IV-A. Table III shows the KCE for the best respective MMUL size for each precision. The table compares the KCC and KCE for three scenarios: unconstrained buffers (BufferOptLevel 9), buffers constrained to the same AIE using buffer location placement (BufferOptLevel 0), and then buffer constrained to the same AIE using our buffer address placement. Unconstrained single AIE buffers show the best performance with the highest average KCE of 94%. However, as mentioned earlier, this is not scalable because the buffers can be allocated in neighboring AIEs by the compiler. It significantly reduces AIE memory utilization as shown in Table II. Since this is the best case performance, we use this as a baseline for further comparison. Buffers constrained to the same AIE using buffer location placement and BufferOptLevel 0 shows an average 16% drop in KCE. The observed drop occurs due to the random fragmentation of buffers across the AIE banks. However, buffers constrained to the same AIE using our custom buffer address placement recover on average 12% of this loss, bringing the average KCE to 90%, while maximizing AIE memory utilization and ensuring scalability.
@@
  • (M, K, N) denotes the final size of the GEMM operation mapped to the AIE array.
@@ V-C Performance of a pack of AIEs
Figure 3 shows the AIE connected as a pack. Each AIE in the pack uses the kernel size from the previous section, optimizing compute to communication ratio $\gamma$, AIE memory utilization, and KCE. We use a pack size of 4 as discussed in Section IV-B. Table IV shows the effect of different placement constraints on the pack. Unconstrained buffers is considered as the baseline with the highest average KCE of 88%. Buffers constrained to the pack using buffer location placement show 14% loss in KCE. Using our custom buffer address placement, we on average recover 12% of this loss and bring the average KCE to 86%.
@@ V-E Comparing with other state-of-the-art designs
Our work is the first work to map GEMM operations on AIE2 devices. Prior work [20, 14, 8, 15] use AIE1. Hence, we only compare relative achieved performance with prior work. Relative comparison is done using Throughput Efficiency (TE) i.e. ratio of achieved throughput to peak throughput. Moreover, different prior works use different precisions - some use int8-int8 [20, 21, 24, 22], int8-int16 [8], int8-int32 [14]. No other prior work comprehensively use all the 4 precisions like GAMA. Hence, we can only compare a part of our results with each prior work. Table VI compares our approach with [8, 20, 24, 22] different works on Versal VC1902 on VCK190 board.
We compare int8 input and int32 output with MAXEVA [14], since that is the only work that writes 32-bit output back to the PL. GAMA achieves 133 TOPs (69%) TE. MaxEVA achieves 77 TOPs (60%) TE. Thus, GAMA extracts 9% more TE. MAXEVA’s array efficiency is limited to 80% since only 80% of the AIEs in the design do matrix multiplication operation (20% are used for addition).
@@
Lastly, we compare GAMA with CHARM [24] and ARIES [22] for int8-int8 precision. Our design outperforms CHARM and ARIES by 53.6% and 39%, respectively, in terms of TE. RSN-XNN [15] only demonstrates performance for fp32. Since AIE2 lacks native support for fp32, a direct comparison between GAMA and RSN-XNN cannot be made.
@@ VI Conclusion
In this work, we introduce the GAMA framework that accelerates GEMM on AMD Versal AIE-ML architecture. Our framework maximizes AIE memory utilization, provides a custom buffer placement algorithm to minimize memory stalls, and a staggered kernel placement to facilitate scaling over the AIE array. GAMA framework provides the highest utilization of peak throughput for Versal compared to all other state-of-the-art frameworks.
Future work will incorporate memory tiles present in the AIE-ML architecture. Maximizing utilization of these memories is critical to improving the overall performance due to Versal’s limited peak off-chip DRAM bandwidth.

### v3

@@ IV-B Chaining the AIEs together
AIE to AIE connectivity: AIEs are connected in series, as shown in Figure 3, to form a pack, allowing larger GEMMs to be implemented with increased parallelism and reuse. There are two PLIOs for input to each AIE in the pack and one PLIO for output from the pack. Reductions are performed inside the pack, and only final results are written back to the PL. This reduces the output PLIOs in the design, since only the last AIE has to write data back to the PL. Cascade, via-switch connection, and buffer are the various methods for partial sum communication from one AIE to the next AIE. CHARM [20], ARIES [22] and RSN-XNN [15] use a cascade interface, while MaxEVA [14] and AMA [8] use buffer. Buffer communication does not stream data from one AIE to another; it uses buffers in the neighboring AIE’s memory. These buffers consume memory. Via-switch and cascade connections stream data from one AIE to another avoiding this memory overhead. Via-switch interface is only 32-bit and using it increases communication latency significantly. AIE architecture has a 512-bit wide cascade interface compared to AIE’s 384-bit. The cascade with the larger bit width and no memory overhead becomes the best candidate for AIE to AIE communication. We use cascade interface for all partial sum communication. Cascade connections make the pack run in a data-flow fashion.
@@
Different kernel types: The pack includes three distinct kernels. The first kernel (used in the leftmost AIE in the pack) takes Matrix A and Matrix B as inputs from the PLIO and output’s partial sum through cascade interface. The second kernel (used in the middle AIEs in the pack) has two input PLIO ports, one input cascade port, and one output cascade port. The third kernel (used in the last AIE in the pack) receives two inputs from PLIO, one input from the cascade port, and one output PLIO port.
@@
AIEs connected in a pack perform a GEMM operation with matrix sizes larger than the matrix sizes for a single AIE. For example, if the kernel sizes are M=48, K=240 and N=48 then the native size of the pack becomes M=48, K=960 and N=48.
@@ IV-C Scaling to the complete AIE array
Scaling hyperparameters and constraints: The scaling in our design is governed by three hyperparameters. Y: Replicates the pack along the Y-axis of the AIE array. G: Defines the size of each pack. X: Replicates the pack along the X-axis of the AIE array. Figure 6 shows the scaling for the case (Y = 2,G = 4,X = 2). PLIOs are shown using different colors that show broadcast and reuse. Matrix A and matrix B are assigned to 16 AIE as shown in Figure 6.
@@
$Y\times X\leq PLIO\_out\$(9)
Kernel placement: Scaling the design to the entire array by satisfying the constraints stated above leads to compilation failure due to PLIO routing congestion. This happens because the third AIE utilizes three distinct PLIOs (two read, one write), leading to congestion in the switch network’s vertical lane when extended across all rows. To avoid this congestion, GAMA uses a zigzag or skewed kernel placement strategy across the array, as shown in Figure 7. We skew the kernel placement by two since skewing by one also leads to congestion, and skewing by three reduces the AIE array utilization. The first two AIEs in each alternate rows of AIEs are not used. The pattern alternates the third AIE’s row position, easing vertical switch lane pressure. This placement successfully avoids congestion, leading to a scalable and compilable design.
@@ V-B Performance of single AIE kernels
At the single AIE level, the two most important aspects are the MMUL API size and the single AIE kernel size (MxKxN). As discussed in Section IV-A, the size of a single AIE kernel depends on factors such as the compute-communication ratio $\gamma$, AIE memory usage and input/output precision.
@@
To calculate the efficiency of the kernel, we measure the kernel compute cycles (KCC) and then take the ratio of theoretical KCC to measured KCC and call it kernel compute efficiency (KCE). We use KCE as a metric to find the best MMUL API size. We sweep all MMUL sizes to find the optimal M, K, and N sizes, as noted in Section IV-A. Table III shows the KCE for the best respective MMUL size for each precision. The table compares the KCC and KCE for three scenarios: unconstrained buffers (BufferOptLevel 9), buffers constrained to the same AIE using buffer location placement (BufferOptLevel 0), and then buffer constrained to the same AIE using our buffer address placement. Unconstrained single AIE buffers show the best performance with the highest average KCE of 94%. However, as mentioned earlier, this is not scalable because the buffers can be allocated in neighboring AIEs by the compiler. It significantly reduces AIE memory utilization as shown in Table II. We use this best-case performance as a baseline for further comparison. Buffers constrained to the same AIE using buffer location placement and BufferOptLevel 0 shows an average 16% drop in KCE. The observed drop occurs due to the random fragmentation of buffers across the AIE banks. However, buffers constrained to the same AIE using our custom buffer address placement recover on average 12% of this loss, bringing the average KCE to 90%, while maximizing AIE memory utilization and ensuring scalability.
@@
  • (M, K, N) denotes the final size of the GEMM.
@@ V-C Performance of a pack of AIEs
Figure 3 shows the AIE connected as a pack. Each AIE in the pack uses the kernel size from the previous section, optimizing compute to communication ratio $\gamma$, AIE memory utilization, and KCE. We use a pack size of 4 as discussed in Section IV-B. Table IV shows the effect of different placement constraints on the pack. Unconstrained buffers is considered as the baseline with the highest average KCE of 88%. Buffers constrained to the pack using buffer location placement show 14% loss in KCE. Our custom buffer placement recovers 12% of the loss, bringing the average KCE to 86%.
@@ V-E Comparing with other state-of-the-art designs
Our work is the first work to map GEMM operations on full AIE-ML array devices. Prior work [20, 14, 8, 15] use AIE. Hence, we only compare the relative achieved performance with prior work. Relative comparison is done using Throughput Efficiency (TE) i.e. ratio of achieved throughput to peak throughput. Moreover, different prior works use different precisions - some use int8-int8 [20, 21, 24, 22], int8-int16 [8], int8-int32 [14]. No other prior work use all the 4 precisions like GAMA. Hence, we can only compare a part of our results with each prior work. Table VI compares our approach with [8, 20, 24, 22] different works on Versal VC1902.
We compare int8 input and int32 output with MAXEVA [14], since that is the only work that writes 32-bit output back to the PL. GAMA achieves 133 TOPs (69%) TE. MaxEVA achieves 77 TOPs (60%) TE. GAMA extracts 9% more TE. MAXEVA’s efficiency is capped at 80% as only 80% of AIEs perform matrix multiplication, with 20% used for addition.
@@
Lastly, we compare GAMA with CHARM [24] and ARIES [22] for int8-int8 precision. Our design outperforms CHARM and ARIES by 53.6% and 39%, respectively, in terms of TE. RSN-XNN [15] only demonstrates performance for fp32. Since AIE-ML lacks native support for fp32, a direct comparison between GAMA and RSN-XNN cannot be made.
@@ VI Conclusion
In this work, we introduce the GAMA framework that accelerates GEMM on AMD Versal AIE-ML architecture. Our framework maximizes AIE memory utilization, provides a custom buffer placement algorithm to minimize memory stalls, and a staggered kernel placement to facilitate scaling over the AIE array. GAMA framework provides the highest utilization of peak throughput for Versal compared to all other state-of-the-art frameworks. Though our improvements focus on AIE-ML, optimizing memory usage and prioritizing scalability can also benefit other accelerator architectures.
Future work will incorporate the memory tiles present in the AIE-ML architecture. Maximizing utilization of these memories is critical to improving the overall performance due to Versal’s limited peak off-chip DRAM bandwidth.